answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability that the dealer will not get blackjack?

The chance of a Blackjack appearing for any player OR the dealer is slightly different depending solely upon how many decks are used.

The approximately correct answer in all cases though is odds of 20-1

That is to say, that one out of every 21 hands is a Blackjack.

Ergo, the probability of the occurrence is approx. 4.8%

The question asked is "What is the probability of the Dealer not getting a blackjack."

Therefore, the correct answer is 100% - 4.8% = 95.2%

If you consider that there is 1 Ace per suit, and 4 tens per suit, then the correct answer is (1/13 * 4/13) = 4/169 or a 2.367% chance that the dealer will get blackjack.

The chance that with an Ace showing that the dealer has a 10 under the Ace, is 4/13 or 30.769%.

With a 10 showing, the chance the dealer has an Ace below the 10 is 1/13 or 7.692%.


The reason you aren't supposed to take insurance, is because of the 30% chance to have it. However if the deck is +6, then even a professional card counter will take insurance.

Probability is the what that something will or will not occur?

Probability is the likelihood that something will occur. If you subtract it from 1, we get the likelihood (or probability) that it will not occur.

If a coin is tossed and rolls heads 6 times, the (empirical) probability of obtaining a head is 6/10 or .6.

1-.6 =.4 is the empirical probability (or likelihood) of not getting a head.

In maths What is Mutually exclusive probability?

Mutually exclusive means they are independent of one another. So, the two events are independent of one another.

Working models of mathematics?

yes ,I knew it , u r loooser and also u have lost ur imaginery sense to make simple model

What is the probability of getting tails and selecting a 3 if you toss a coin and randomly select a number?

There are infinitely many numbers and so the probability of the second event is 0. As a result the overall probability is 0.

How do you get a black card?

You get a black card by drawing it.

Example of theoretical probability?

Roll a fair cubic die. The theoretical probability of it ending up with any particular face on top is 1/6

4 coins tossed together. What is the probability of getting 2 tails?

The probability of it is 37.5%. Since there are 16 possible outcomes. calculated by 2 to the power of 4. only 6 gives you the prefered results then it's a simple 6:16 ratio

What is a phlidamide?

somebody with no arms or legs

If a student is picked at random what is the probability that heshe received an A on hiher final?

If a student is picked at random what is the probability that he/she received an A on his/her fina?

If a rifleman hits target 80 percent of the time If he fires four shots what is the probability he will hit the target exactly twice and what is the probability he will hit the target at least twice?

Probabilities are calculated by the binomial distribution. Ans: 0.1536 and 0.9728

Discussion:

Prob of exactly two times = 4!/(2! x 2! ) x (0.8)^2 x (0.2)^2 = 0.1536

Prob of at least twice = hits target 2 times + hits target 3 times + hits target 4 times

Prob of at least twice = 1 - Prob of no hits - Prob of one hit

Prob of one hit = 4!/(3! x 1!) x (0.8)^1 x (0.8)^1 x (0.2)^3 = 0.0256

Prob of no hits= 4!/(0! x 4!) x (0.8)^0 x (0.2)^4 = (0.2)^4 = 0.0016

Prob of 2 or more hits = 1-0.0256-0.0016 = 0.9728

Note: You can calculate these values using Excel, where Prob of 2 hits = binom(2,4,0.8,false)

and Prob of 2 or more hits = 1 - binom(1,4,0.8,false) - binom(0,4,0.8,false)

or:

Prob of 2 > hits = 1- binom(1,4,0.8,true) , as false is requesting the PMF value and true is requesting CDF value. See help in Excel for further explanation on this function.

Also, there might be an issue of independent events in this problem, in that the probability is given as a constant. If a rifleman missed three times, do you think he would learn and might do better than his average on the fourth shot?

Name two characters who were seeking in great carbuncle Include their motivation for seraching for the gem and each of their outcomes?

The young Couple to light up there house in the winter evenings.

The Poet To just look at its beauty to inspire him in writing poems.

If you have a 52 card deck and deal 48 cards what is the probability or percentage that the Jack of Spades is in the remaining 4 cards?

If you have a 52 card deck and deal 48 cards, the probability that the Jack of Spades

is in the remaining 4 cards is the same as in the case of dealing 4 cards from the 52

card deck, the probability the Jack of Spades is in those 4 cards.

To solve the problem you can see there are 4 possible ways that the Jack of Spades

can come up in the 4 deal-led cards.

I.- It can come in the first card. P1= 1/52.

II.- Come in the second card. P2= 51/52 (1/51).

III.- Come in the third card. P3= 51/52 (50/51) (1/50).

IV.- Or in the fourth card. P4= 51/52 (50/51) (49/50) (1/49).

The probability looked for is P = P1+P2+P3+P4 = 1/13

A coin is tossed five times find the probability of getting exactly three heads?

Since a coin has two sides and it was tossed 5 times, there are 32 possible combinations of results. The probability of getting heads three times in 5 tries is 10/32. This is 5/16.

Trending Questions
You are dealt one card from a standard 52 card deck Find the probability of being dealt a club and a spade? What does sample mean in an entire population? A multiple choice test has 8 questions that has 3 possible answers only one is correct If Judy who forgot to study for the test guesses on all questions what is the probability that she gets 3 right? When you label a ray with 2 letters can you write leters both ways? What is aceptable drawing accuracy standard? How many red kings are there in a standard deck of cards? What is the probabilitty of flipping a coin three times and getting taills? Does the standard normal distribution is unimodal? What is the probability of rolling a sum which is a multiple of five? How many ways can the letters in the word mathematica be arranged? What is the probability getting a thirteen by rolling two number cubes? What is the probability that the dice lands as a number less than 3? Why is it hard to calculate theoretical probability for bowling? What is the probability of rolling a number that is divisible by 2 on a six sided die? Why is probability usually described with fractions? How many queens ace or kings are there in a deck of cards? What is the probability of rolling doubles when two dice are rolled? What is the sample whose characteristics are the same as those of the population it is intended to reflect? The likelihood that an event will occur like tossing a coin? An area of a normal probability distribution represents?